OpenStudy (tmyrci):

(X^2+3y) (3y+7) x=1 y=1

alones (alones):

Okay for this I got 40. You put forx=1 and for 1 you put also 1 like this \[\left(1^2+3\left(1\right)\right)\left(3\left(1\right)+7\right)\] \(\ 40\)